### v2

# Graph Attention with Random Rewiring

Abstract
Graph Neural Networks (GNNs) have become fundamental in graph-structured deep learning. Key paradigms of modern GNNs include message passing, graph rewiring, and Graph Transformers. This paper introduces Graph-Rewiring Attention with Stochastic Structures (GRASS), a novel GNN architecture that combines the advantages of these three paradigms. GRASS rewires the input graph by superimposing a random regular graph, enhancing long-range information propagation while preserving structural features of the input graph. It also employs a unique additive attention mechanism tailored for graph-structured data, providing a graph inductive bias while remaining computationally efficient. Our empirical evaluations demonstrate that GRASS achieves state-of-the-art performance on multiple benchmark datasets, confirming its practical efficacy.
1 Introduction
Graph neural networks (GNNs) have emerged as a key innovation in machine learning for tasks involving complex relational data represented as graphs [67, 72]. Central to the advancement of GNNs are message passing neural networks (MPNNs), exemplified by the Graph Convolutional Network (GCN) [39]. By aligning computation with the structure of the graph, MPNNs offer a strong inductive bias for graph-structured data. However, despite recent advances, limitations persist in challenges such as underreaching, which occurs when distant nodes cannot communicate [1], oversmoothing, which occurs when node representations become indistinguishable [52, 15], and oversquashing, which occurs when task-relevant information is excessively compressed [7, 64]. Graph rewiring techniques, aimed at modifying the graph structure to enhance information flow, have shown promise in overcoming these limitations [2]. However, it comes with the risk of distorting the structure of the graph, leading to loss or misinterpretation of relational information [20].
Building GNNs with Transformers [65] presents another promising solution to the above-listed problems. However, many Graph Transformers (GTs) face scalability challenges as a result of their reliance on quadratic-time Transformers [56]. This makes their applicability to large graphs computationally challenging. To address these issues, the General, Powerful, Scalable (GPS) Graph Transformer [56] combines linear-time Transformers with MPNNs to offer a method with $O(|V|+|E|)$ time complexity, where $|V|$ denotes the number of vertices and $|E|$ the number of edges. Exphormer [61], which generalizes sparse Transformers [76] to graphs, represents another promising approach. All these architectures, however, inherit their attention mechanisms from Transformers that are intended for sequences, which are often not optimal for graph-structured data.
Our Contribution.
We aim to design a GNN that encompasses the graph inductive bias offered by message passing, the improved information flow achieved through graph rewiring, and the representation power of the attention mechanism. Towards this end, we propose Graph-Rewiring Attention with Stochastic Structures (GRASS), a novel GNN inspired by these three paradigms. GRASS rewires the input graph by superimposing a random regular graph. At each layer, it applies an attentive node aggregator with attention weights computed from edge representations, and an MLP edge aggregator that updates edge representations to reflect node relationships.
We demonstrate 1) the intuitive and empirical effectiveness of combining random rewiring with message passing, 2) the advantage of a graph-tailored additive attention mechanism over Transformer-style dot-product attention, and 3) the link between randomly rewired message passing and sparse GTs through two interpretations of GRASS. In addition, we show that GRASS achieves state-of-the-art performance on multiple datasets.
In Section 2, we briefly review some related work, highlighting their strengths and weaknesses. In Section 3, we discuss the design principles, architecture, and interpretations of GRASS, our novel GNN. We demonstrate the performance of GRASS in Section 4. We finish the paper with some concluding remarks in Section 5.
2 Related Work
In this section, we briefly summarize some of the prior work related to the main concepts of GRASS.
Graph Rewiring.
Graph rewiring involves the strategic modification of graph structures by adding, removing, or reallocating edges. Techniques include deterministic adjustment of the graph structure based on spectral properties [29, 64, 2, 20] and probabilistic rewiring of the graph based on learned distributions [54]. DRew [32] innovatively introduces layer-dependent rewiring, demonstrating outstanding performance on tasks that depend on long-range relationships. While graph rewiring enhances MPNNs by improving graph connectivity, and thus information propagation, they risk compromising informative structural properties of graphs, degrading its inductive bias [20].
Graph Attention and Graph Transformers.
Graph Attention Network (GAT) [68] and its successors [9, 78, 70, 11] have leveraged similar attention-based approaches to achieve remarkable results [9, 25]. Despite their success, these MPNNs are intrinsically limited by the structure of the input graph, leading to persistent challenges such as underreaching [1], oversmoothing [52, 15], and oversquashing [7, 64]. At a theoretical level, the expressivity of MPNNs is often no more powerful than the Weisfeiler-Lehman graph isomorphism test [73, 60, 41].
Graph Transformers [75, 22, 40, 74, 48, 36] employ the Transformer architecture [65] to transcend these limitations of MPNNs. Instead of relying on message passing within neighborhoods of the input graph, GTs facilitate simultaneous attention between all nodes, unrestricted by the structure of the graph. However, its $O(|V|^{2})$ time complexity becomes prohibitive for large graphs. Linear and sparse Transformers, exemplified by Performer [18] and BigBird [76], respectively, achieve linear runtime with respect to the number of tokens. However, integrating a graph inductive bias into these Transformers can be challenging, since they are designed for Euclidean data [47].
The General, Powerful, Scalable (GPS) Graph Transformer [56] represents a hybrid approach, merging MPNNs’ inductive bias with the global perspective of Linear-Time Transformers. Exphormer [61] presents a natural generalization of BigBird to graphs by incorporating local, global, and random attention. It employs random regular expander graphs as an alternative to BigBird’s random attention patterns on sequential data, and is analogous to the random rewiring mechanism of GRASS. However, these models, along with prior GTs, focus on adapting Transformer-style dot-product attention to graph-structured data, rather than investigating graph-specialized alternatives. Their performance does not exceed that of more tailored attention mechanisms, such as GRIT [47], on a variety of tasks.
Graph Encoding.
By integrating structural information directly into nodes and edges, graph encoding significantly improves the expressivity of GNNs, especially for GTs [25]. Techniques such as graph Laplacian positional encoding (LapPE) [25] and random walk structural encoding (RWSE) [23] have been instrumental in this regard. The Graph Inductive Bias Transformer (GRIT) [47] introduces a novel encoding strategy—relative random walk probabilities (RRWP)—which is particularly powerful in capturing structural properties [47]. However, like its GT predecessors, GRIT suffers from $O(|V|^{2})$ complexity, limiting its applicability to large graphs.
In summary, while existing methodologies have significantly advanced the field of graph neural networks, they are often constrained by limited expressivity within the MPNN framework, computational inefficiencies, or challenges in incorporating a graph inductive bias. GRASS emerges as a novel solution that addresses these challenges through the combination of random rewiring and a graph-tailored attention mechanism, offering a competitive solution to graph-structured learning.
3 Methods
In this section, we introduce the design of GRASS. We begin by examining the desirable qualities of a GNN, which guide our architectural design. Subsequently, we introduce the components of GRASS by following the order of data processing in our model, and also describe the role of each component.
Design Goals.
We center our design around what we consider to be the key characteristics of an effective GNN. We will focus on the processing of nodes (N1–N3) and edges (E1–E2), as well as the scalability (S1) of the proposed model.
  1. N1.
    Permutation Equivariance. Unlike tokens in a sentence or pixels in an image, nodes in a graph are unordered, and therefore the model should be permutation equivariant by construction. Since reordering the nodes of a graph does not change the graph, permuting the nodes of the input graph of a GNN layer should only result in the same permutation of its output [67]. Formally, let $f(\mathbf{X},\mathbf{E},\mathbf{A})$ be the function computed by a layer, where $\mathbf{X}:\mathbb{R}^{|V|\times n_{\text{node}}}$ represents node features with $n_{\text{node}}$ dimensions, $\mathbf{E}:\mathbb{R}^{|V|\times|V|\times n_{\text{edge}}}$ represents edge features with $n_{\text{edge}}$ dimensions, and $\mathbf{A}:\{0,1\}^{|V|\times|V|}$ represents the adjacency matrix (edge weights are considered scalar-valued edge features). If the layer is permutation equivariant and $(\mathbf{X}_{\text{out}},\mathbf{E}_{\text{out}})=f(\mathbf{X}_{\text{in}},\mathbf{E}_{\text{in}},\mathbf{A})$, then $(\mathbf{PX}_{\text{out}},\mathbf{PE}_{\text{out}}\mathbf{P}^{\top})=f(\mathbf{PX}_{\text{in}},\mathbf{PE}_{\text{in}}\mathbf{P}^{\top},\mathbf{PAP}^{\top})$ for an arbitrary permutation matrix $\mathbf{P}$.
  2. N2.
    Effective Communication – The model should facilitate long-range communication between nodes. Numerous real-world tasks require the GNN to capture interactions between distant nodes [24]. However, MPNN layers, which propagate information locally, frequently fail in this regard [47]. A major challenge is underreaching, where an MPNN with $l$ layers is incapable of supporting communication between two nodes $x_{i},x_{j}$ with distance $\delta(x_{i},x_{j})>l$ [1]. Another challenge is oversquashing, where the structure of the graph forces information from a large set of nodes to squeeze through a small set of nodes to reach its target [64]. A node with a constant-size feature vector may need to relay information from exponentially many nodes (with respect to model depth), leading to excessive compression of messages in deep MPNNs [1]. Addressing these issues is crucial for effectively capturing long-distance relationships.
  3. N3.
    Selective Aggregation – The model should only aggregate information from relevant nodes and edges. MPNN layers commonly update node representations by unconditionally summing or averaging messages from neighboring nodes and edges [67]. In deep models, this can lead to oversmoothing, where the representation of nodes from different classes becomes too similar in the process of repeated aggregation [15]. MPNN like GCN [39] have been shown to behave as low-pass filters in feature vectors [52]. To address this issue, nodes should only aggregate information from relevant neighbors, instead of doing so unconditionally, in order to maintain distinguishability of node representations when required by the task.
  4. E1.
    Relationship Representation – The model should effectively represent the relationships between nodes with edges. Edges in graph-structured data often convey meaningful information about the relationship between the nodes it connects [30]. However, many contemporary GNNs do not use edge features [33, 73, 48], only support scalar-valued edge features (edge weights) [39, 21], or never update edge features across layers [74, 61]. In addition to the semantic relationship represented by edge features of the input graph, structural relationship can be represented by edge encodings added by the model [56]. To inform the model about the relationship between nodes, edge representations should incorporate information from both edge features and edge encodings, and be meaningfully utilized by layers.
  5. E2.
    Directionality Preservation – The model should preserve and utilize information carried by edge directions. Many graphs representing real-world relationships are inherently directed [58]. Although edge directionality has been shown to carry important information for various tasks, many GNN variants require undirected graphs as input, to prevent edge directions from restricting information flow [58]. It would be beneficial for the model’s expressivity if edge directionality information could be preserved without severely limiting communication.
  6. S1.
    Efficient Computation – The model should have $O(|V|+|E|)$ time complexity for its non-precomputable operations. The number of edges $|E|$ in a connected directed graph with $|V|$ vertices may vary from $|V|$ to $|V|(|V|-1)$. However, in practical scenarios, the number of edges typically falls somewhere between these extremes. GNNs with $O(|V|)$ complexity can be incapable of processing every edge, while those with $O(|V|^{2})$ complexity struggle to scale to large-and-sparse graphs where $|E|\in O(|V|)$. Meanwhile, a model with $O(|V|+|E|)$ complexity can be efficient on both dense and sparse graphs. While this is tempting, we notice that computing certain graph properties, such as detecting triangles [71] and finding unweighted all-pair shortest paths [14], cannot be done in $O(|V|+|E|)$ time, which prohibits a model with $O(|V|+|E|)$ time complexity from completing tasks that require these properties. As a compromise, we allow the model to preprocess data without this restriction to capture critical structural properties, as long as the model’s non-precomputable steps remain $O(|V|+|E|)$ time. Precomputable operations only need to be performed once per dataset, regardless of the number of training sessions and epochs, thereby reducing their impact on overall efficiency.
The Structure of GRASS.
The high-level structure of GRASS is illustrated in Figure 2. Prior to training, GRASS precomputes the degree matrix and random walk probabilities of each graph in the dataset. At each training iteration, GRASS randomly rewires the input graph, applies node and edge encodings, and passes the graph through multiple attention layers, producing an output graph with the same structure. For tasks that require a graph-level representation, such as graph regression and graph classification, pooling is performed on the output graph to obtain a single output vector.
3.1 Graph Encoding
Extracting structural information plays an important role in graph-structured learning and is crucial for Relationship Representation. To this end, GRASS applies relative random walk probabilities (RRWP) encoding [47] and degree encoding [74] to represent structural relationships.
RRWP Encoding.
RRWP encoding has been shown to be very expressive both theoretically and practically [47], serving as a major source of structural information for the model. To calculate random walk probabilities, we first obtain the transition matrix $\mathbf{T}$, where $\mathbf{T}_{i,j}$ represents the probability of moving from node $i$ to node $j$ in a random walk step. It is defined as $\mathbf{T}=\mathbf{D}^{-1}\mathbf{A}:[0,1]^{|V|\times|V|}$, where $\mathbf{A}\in\{0,1\}^{|V|\times|V|}$ is the adjacency matrix of the input graph $G$, and $\mathbf{D}\in\mathbb{N}^{|V|\times|V|}$ is its degree matrix. The powers of $\mathbf{T}$ are stacked to form the RRWP tensor $\mathbf{P}$, with $\mathbf{P}_{h,i,j}$ representing the probability that a random walker who starts at node $i$ lands at node $j$ at the $h$-th step. Formally, $\mathbf{P}=[\mathbf{T},\mathbf{T}^{2},...,\mathbf{T}^{k}]:[0,1]^{k\times|V|\times|V|}$, where $k$ is the number of random walk steps. The diagonal elements $\mathbf{P}_{:,i,i}$ where $i\in V_{G}$ are used as node encodings, similarly to RWSE [23]. The rest are used as edge encodings when the corresponding edge is present in the rewired graph $H$. All node encodings undergo Batch Normalization (BN) [37] to improve their distribution, while batch noise also acts as a regularizer that increases the model’s resistance to noise in the random walk probabilities produced by structural noise in the input graph. Here, $n$ denotes the dimensionality of hidden layers.
$\mathbf{x}_{i}^{\text{RW}}$ $=\mathbf{W}_{\text{node-enc}}\text{BN}(\mathbf{P}_{:,i,i}):\mathbb{R}^{n}$ (1)
$\mathbf{e}_{i,j}^{\text{RW}}$ $=\mathbf{W}_{\text{edge-enc}}\text{BN}(\mathbf{P}_{:,i,j}):\mathbb{R}^{n}$ (2)
As required by Efficient Computation, operations slower than $O(|V|+|E|)$ time must be precomputable. $\mathbf{P}$ only needs to be computed once per dataset because it depends only on the adjacency matrix $\mathbf{A}$ of the input graph. To further improve performance through pipelining, the CPU looks up $\mathbf{P}_{:,i,i}$ for each $i\in|V_{G}|$ and $\mathbf{P}_{:,i,j}$ for each edge $(i,j)\in E_{G}$ to hide latency and conserve VRAM.
Degree Encoding.
Degree encoding is another simple yet powerful way to represent structural information [74]. For datasets in which node degrees can take a small number of values, the model could learn a degree embedding tensor $\mathbf{E}:\mathbb{R}^{D^{+}\times D^{-}\times n}$, where $D^{+}$ and $D^{-}$ denote the maximum out- and in-degree of the dataset, while $\text{d}^{+}(i)$ and $\text{d}^{-}(i)$ represent the out- and in-degree of node $i$.
$\mathbf{x}_{i}^{\text{deg}}=\mathbf{E}_{\text{d}^{+}(i),\text{d}^{-}(i),:}:\mathbb{R}^{n}$ (3)
[…]
$\mathbf{x}_{i}^{\text{deg}}=\mathbf{W}_{\text{deg}}\text{BN}(\text{d}^{+}(i)\,\|\,\text{d}^{-}(i)):\mathbb{R}^{n}$ (4)
Applying Encodings.
Before entering attention layers, RRWP encodings are added to both node features and edge features (including those of edges added by random rewiring), while degree encodings are only added to node features.
$\mathbf{x}_{i}^{0}$ $=\mathbf{x}^{\text{in}}_{i}+\mathbf{x}^{\text{RW}}_{i}+\mathbf{x}^{\text{deg}}_{i}:\mathbb{R}^{n}$ (5)
$\mathbf{e}_{i,j}^{0}$ $=\mathbf{e}^{\text{in}}_{i,j}+\mathbf{e}^{\text{RW}}_{i,j}:\mathbb{R}^{n}$ (6)
3.2 Random Rewiring
To address the issues of underreaching and oversquashing for Effective Communication, GRASS rewires the input graph by superimposing a random regular graph. While deterministic rewiring and non-regular graphs may also improve connectivity, we present potential advantages of using random regular graphs. We also provide details on random regular graph generation and input graph rewiring.
Effects on Diameter.
The diameter of a graph upper-bounds the distance between two nodes, and thus the number of layers for an MPNN to propagate information between them [1]. Superimposing a random regular graph onto the input graph can drastically decrease its diameter. The least integer $d$ that satisfies $(r-1)^{d-1}\geq(2+\varepsilon)r|V|\log|V|$ is the upper bound of the diameter of almost every random $r$-regular graph with $|V|$ nodes, where $r\geq 3$ and $\varepsilon>0$ [8]. Since adding edges to a graph never increases its diameter, the diameter $d$ of the rewired graph is asymptotically upper-bounded by $d\in O(\log_{r}|V|)$ when $r\geq 3$. Subsequently, all nodes would be able to communicate with each other given $O(\log_{r}|V|)$ message passing layers, which could significantly reduce the risk of underreaching on large graphs. In addition, the diameter of a graph is a trivial upper bound of its effective resistance [26], which has been shown to be positively associated with oversquashing [7]. Intuitively, it upper bounds the “length” of the bottleneck through which messages are passed.
Effects on Internally Disjoint Paths.
Since oversquashing can be attributed to squeezing too many messages through the fixed-size feature vector of a node [1], increasing the number of internally disjoint paths between two nodes may reduce oversquashing by allowing information to propagate through more nodes in parallel. Intuitively, it increases the “width” of the bottleneck. A random $r$-regular graph with $r\geq 2$ almost certainly has a vertex connectivity of $r$ as $|V|\rightarrow\infty$ [27]. Menger’s Theorem then lower-bounds the number of internally disjoint paths by a graph’s vertex connectivity [31]. This provides another perspective on how the random regular graph may reduce oversquashing.
Effects on Spectral Gap.
Oversquashing has been shown to decrease as the spectral gap of a graph increases, which is defined as $\lambda_{1}$, the smallest positive eigenvalue of the graph’s Laplacian matrix [38]. It has been proven that a random $r$-regular graph sampled uniformly from the set of all $r$-regular graphs with $|V|$ nodes almost certainly has $\mu<2\sqrt{r-1}+1$ as $|V|\rightarrow\infty$, where $\mu$ is the largest absolute value of nontrivial eigenvalues of its adjacency matrix [53]. Since the graph is $r$-regular, its $i$-th adjacency matrix eigenvalue $\mu_{i}$ and $i$-th Laplacian matrix eigenvalue $\lambda_{i}$ satisfy $\lambda_{i}=r-\mu_{i}$ [46], lower-bounding the spectral gap with $\lambda_{1}>r-2\sqrt{r-1}-1$. Therefore, the superimposed random regular graph may alleviate oversquashing by increasing the spectral gap of the rewired graph.
Generating Random Regular Graphs.
We generate random regular graphs with the Permutation Model [53] that we describe here and with pseudocode in Appendix A. For a given positive and even parameter $r\geq 2$, and for the input graph $G=(V_{G},E_{G})$, we randomly generate a corresponding $r$-regular graph by independently and uniformly sampling $\frac{r}{2}$ random permutations $\sigma_{1},\sigma_{2},...,\sigma_{\frac{r}{2}}$ from $S_{|V_{G}|}$, the symmetric group defined over the nodes of the graph $G$. Using these random permutations, we construct a random pseudograph $\tilde{R}=(V_{G},E_{\tilde{R}})$, where the edge set $E_{\tilde{R}}$ of the graph $\tilde{R}$ is
$E_{\tilde{R}}=\bigsqcup_{i\in V_{G},j\in[\frac{r}{2}]}\{\{i,\sigma_{j}(i)\}\}$ (7)
Here, $\sqcup$ denotes the disjoint union of sets. The resulting graph $\tilde{R}$ is a random regular pseudograph, and the probability that $\tilde{R}$ is any regular pseudograph with $|V_{G}|$ nodes and degree $r$ is equal [8]. Being a pseudograph, $\tilde{R}$ might not be simple: it might contain self-loops and multi-edges. Even when $|V_{G}|$ is large, the probability that $\tilde{R}$ is simple—that it does not contain self-loops or multi-edges—would not be prohibitively small. In particular, it has an asymptotically tight lower bound [8]:
$\lim_{|V_{G}|\rightarrow\infty}\text{Pr}[\tilde{R}\text{ is simple}]=e^{-\frac{r^{2}}{2}-r}$ (8)
Therefore, if we re-generate $\tilde{R}$ when it’s not simple, the expected number of trials required for us to obtain a simple $\tilde{R}$ is upper-bounded by $e^{\frac{r^{2}}{2}+r}$, which can be kept low by keeping $r$ low. In practice, GRASS would like the superimposed graph to be simple, in order to avoid unwanted redundancy in the attention mechanism. In particular, self-loops result in a duplicate path of message passing from a node to itself, while multi-edges result in multiple identical messages. Meanwhile, the regularity of the graph is desired but not strictly required. Therefore, when $\tilde{R}$ is not simple, we simply remove self-loops and multi-edges from $\tilde{R}$ to obtain $R$, which is always simple but not necessarily regular.
Rewiring the Input Graph.
Figure 1: Visualization of the proposed random rewiring mechanism. Solid lines denote existing edges of the input graph, and dashed lines denote added edges. (a) The input graph $G$. (b, c) Two among all possible instances of the randomly rewired graph $H$ with $r=2$.
To rewire the input graph $G$, GRASS superimposes the edges of $R$ onto $G$, producing a new graph $H=(V_{G},E_{G}\sqcup E_{R})$ that is used as input for subsequent stages of the model. Since it is possible that $E_{G}\cap E_{R}\neq\emptyset$, there might be multi-edges in $H$, and in these cases, $H$ is not a simple graph. GRASS does not remove these multi-edges to avoid biasing the distribution of the superimposed random regular graph.
The added edges $E_{R}$ are given a distinct embedding to make them distinguishable from the existing edges $E_{G}$. This aids Selective Aggregation, as it enables a node to select between its neighbors and a random node. Edge encodings are also given to the added edges, which are instrumental in Relationship Representation. Although an added edge $(i,j)\in E_{R}$ lacks edge features that represent semantic relationships in the input graph, the structural relationship between nodes $i$ and $j$ is represented by the random walk probabilities $\mathbf{P}_{:,i,j}$ given to that edge as its RRWP encoding.
3.3 Attention Mechanism
Many GTs emulate the structure of Transformers designed for Euclidean data [22, 40, 74, 36, 61]. Meanwhile, GRASS uses attentive node aggregators with attention scores computed by MLP edge aggregators, which is a more tailored attention mechanism for graph-structured data. Figure 3 illustrates its structure in detail. The proposed attention mechanism is unique in the way it uses edge representations as the medium of attention weights. To satisfy the needs of Relationship Representation, edge representations must be updated alongside node representations. Unlike a node, whose neighborhood is a set of nodes, a directed edge always has one head and one tail which is an ordered pair. An undirected edge can be represented by two directed edges with opposite directions, which allows us to generalize this observation to undirected graphs. With an ordered input, an edge can use an MLP as its aggregator, offering better expressivity by MLP’s universal approximation [35] while preserving Permutation Equivariance. Assuming Relationship Representation is satisfied, edge features would already represent node relationships, which we can use as attention weights $\mathbf{a}_{i,j}$ after applying a linear layer and taking element-wise softmax [10] to obtain a probability distribution.
The proposed attention mechanism is defined as follows, where $\odot$ denotes the Hadamard product, $\hat{\text{d}}^{-}$ denotes in-degree after random rewiring, and $\mathcal{N}^{-}$ denotes in-neighbors.
$\mathbf{s}_{i,j}^{l}$ $=\text{dropout}\left(\hat{\text{d}}^{-}(j)\exp(\mathbf{W}^{l}_{\text{attn}\leftarrow\text{edge}}\mathbf{e}_{i,j}^{l-1})\right):{\mathbb{R}^{+}}^{n}$ (9)
$\mathbf{a}_{i,j}^{l}$ $=\frac{\mathbf{s}_{h,j}^{l}}{\sum_{h\in\mathcal{N}^{-}(j)}\mathbf{s}_{h,j}^{l}+\varepsilon}:{\mathbb{R}^{+}}^{n}$ (10)
$\tilde{\mathbf{x}}_{j}^{l}$ $=\mathbf{W}^{l}_{\text{tail}\leftarrow\text{tail}}\mathbf{x}_{j}^{l-1}+\sum_{i\in\mathcal{N}^{-}(j)}\mathbf{a}_{i,j}^{l}\odot\mathbf{(}\mathbf{W}^{l}_{\text{tail}\leftarrow\text{head}}\mathbf{x}_{i}^{l-1}+\mathbf{W}^{l}_{\text{tail}\leftarrow\text{edge}}\mathbf{e}_{i,j}^{l-1}):\mathbb{R}^{n}$ (11)
$\tilde{\mathbf{e}}_{i,j}^{l}$ $=\mathbf{W}^{l}_{\text{edge}\leftarrow\text{edge}}\mathbf{e}_{i,j}^{l-1}+\mathbf{W}^{l}_{\text{edge}\leftarrow\text{head}}\mathbf{x}_{i}^{l-1}+\mathbf{W}^{l}_{\text{edge}\leftarrow\text{tail}}\mathbf{x}_{j}^{l-1}:\mathbb{R}^{n}$ (12)
In calculating attention weights, GRASS scales the input to the softmax operation by the logarithm of the number of keys at each node, which is equal to the in-degree of that node after random rewiring. This has been shown to reduce the impact of the number of keys on the entropy of the distribution of attention weights when there is a large number of keys [17], which supports Selective Aggregation. Intuitively, this allows a node to concentrate its attention weight on relevant in-neighbors regardless of the number of in-neighbors it has. Since the number of keys a node attends to is equal to its in-degree, this approach is also analogous to degree scaling [19], which is widely used in MPNNs to incorporate degree information into message passing.
In addition, GRASS uses DropKey [42] as a regularizer, which, in the context of graph attention, is effective in reducing its dependence on the presence of particular edges in the graph. This improves the robustness of GRASS against structural noise and thus facilitates Selective Aggregation. DropKey can be seen as the random removal of edges because it randomly masks the attention matrix prior to the softmax operation. Thus, it elegantly complements random rewiring, which only adds edges. This equivalence could not be established for Attention Dropout [77] despite its popularity in Transformers.
Like Transformers, the output of the attention mechanism is passed through an FFN. Here, $\phi$ denotes a ReLU-like [50, 55, 49] nonlinear activation function.
$\hat{\mathbf{x}}_{i}^{l}$ $=\mathbf{W}^{l}_{\text{node-out}}\phi(\tilde{\mathbf{x}}_{i}^{l}+\mathbf{b}^{l}_{\text{node-act}})+\mathbf{b}^{l}_{\text{node-out}}:\mathbb{R}^{n}$ (13)
$\hat{\mathbf{e}}_{j,i}^{l}$ $=\mathbf{W}^{l}_{\text{edge-out}}\phi(\tilde{\mathbf{e}}_{i,j}^{l}+\mathbf{b}^{l}_{\text{edge-act}})+\mathbf{b}^{l}_{\text{edge-out}}:\mathbb{R}^{n}$ (14)
Normalization and Residual Connection.
PN-V is a variant of Power Normalization [59], which normalizes the root mean square of the batch and is more stable than BN on large-variance inputs. BN has been shown to work poorly in GNNs [13], which would be made worse by the variance introduced with random rewiring. PN-V is used by GRASS to address this problem, allowing it to achieve adequate stability during training while still enjoying normalized batch distributions [59].
The residual connection [34] is scaled by $\alpha$, a coefficient calculated according to DeepNorm [69], which stabilizes the training of deep residual networks. This allows us to use post-normalization, which has been shown to increase expressivity while avoiding the problem of exploding model update [45, 69]. This is essential for us to validate the robustness of GRASS against oversquashing and oversmoothing, which deep GNNs are more vulnerable to, by allowing us to train a deeper model.
[…]
Edge Flipping.
While the proposed attention mechanism naturally achieves Directionality Preservation by aggregating information in the same direction as edges, it can severely restrict the flow of information, putting it in conflict with Effective Communication. As a solution, the direction of each edge is switched from one layer to the next: in odd-numbered layers, the directions match those of the edges in the rewired graph $H$, whereas in even-numbered layers, they are reversed. This enables the model to propagate information in both directions even when the input graph is directed, improving its expressivity [58]. We empirically find edge flipping to be beneficial on undirected graphs (where an undirected edge is represented as a pair of directed edges with opposite directions) as well.
Graph Pooling.
[…]
$\mathbf{y}=\;\sum_{i\in V_{G}}\mathbf{x}_{i}^{L}\;\left\|\;\sum_{(i,j)\in E_{G}}\mathbf{e}_{i,j}^{L}\right.\;\left\|\;\sum_{(i,j)\in E_{R}}\mathbf{e}_{i,j}^{L}\right.\;:\mathbb{R}^{3n}$ (17)
3.4 Interpretations of GRASS
A Message Passing Perspective.
GRASS is an MPNN on a noisy graph. In an MPNN, information is propagated along the edges of the input graph, defined by its adjacency matrix [67]. GRASS can be seen as an MPNN that injects additive and multiplicative noise into the adjacency matrix, through random rewiring and DropKey, respectively. The adjacency matrix $\mathbf{A}_{M}$ followed by message passing is given by
$\mathbf{A}_{M}=(\mathbf{A}_{G}+\mathbf{A}_{R})\cdot\mathbf{A}_{D}$ (18)
where $\mathbf{A}_{G}$ is the adjacency matrix of the input graph $G$, $\mathbf{A}_{R}$ is that of the superimposed random regular graph $R$, $\mathbf{A}_{D}$ is a random attention mask sampled by DropKey (which can also be seen as the adjacency matrix of a random graph), $+$ denotes element-wise OR, and $\cdot$ denotes element-wise AND.
Noise injection is well known as an effective regularizer [51], and for graph-structured data, the random removal of edges has demonstrated regularization effects [57]. By adding noise to the adjacency matrix that guides the propagation of information in message passing, GRASS is encouraged to develop robustness against structural noise, which is prevalent in graph-structured data [66].
A Graph Transformer Perspective.
GRASS is a sparse Graph Transformer. Graph Transformers allow each node to aggregate information from other nodes through graph attention mechanisms, with a general definition [67] being
$\mathbf{x}^{\prime}_{j}=\phi\left(\mathbf{x}_{j},\bigoplus_{i\in\mathcal{N}(j)}a(\mathbf{x}_{i},\mathbf{x}_{j})\psi(\mathbf{x}_{i})\right)$ (19)
where $\phi$ and $\psi$ are neural networks, $a$ is an attention weight function, and $\bigoplus$ is a permutation-invariant aggregator. Many GTs compute attention weights using scaled dot-product attention [65], with node features as keys and queries. However, we observe that edge features in GRASS, which are updated by aggregating information from its head and tail nodes with an MLP, could be used to directly compute attention weights as a form of additive attention [5]. Relationship Representation would then be critical for the attention weights to be meaningful, which GRASS satisfies through expressive edge encodings and the deep processing of edge features.
Many GTs achieve sparsity by integrating [56] or generalizing [61] BigBird’s sparse dot-product attention. Meanwhile, GRASS achieves sparsity in a graph-native way: attention is always local, so nonadjacent nodes in the rewired graph would naturally never attend to each other. Seeing GRASS as a Transformer, its attention mask would be $\mathbf{A}_{M}$ as defined in Equation 18. Moreover, there might be little computational benefit in using sparse dot-product attention on graphs-structured data, since its efficient implementation relies on efficient sparse matrix multiplication [76], which is challenging [28, 12] due to the lack of exploitable sparsity patterns in many real-world graphs.
4 Experiments
Benchmarking GRASS.
To measure the performance of GRASS, we train and evaluate it on five of the GNN Benchmark Datasets [25]: ZINC, MNIST, CIFAR10, CLUSTER, and PATTERN, as well as three of the Long Range Graph Benchmark (LRGB) [24] datasets: Peptides-func, Peptides-struct, and PascalVOC-SP. Following the experimental setup of Rampášek et al. [56], we configure GRASS to around 100k parameters for MNIST and CIFAR10, and 500k parameters for all other datasets. Additional information on the datasets can be found in Appendix D. In all experiments, we use the Lion optimizer [16] and the 1cycle learning rate schedule [62].
As shown in Tables 1 and 2, GRASS ranks first in ZINC, MNIST, CIFAR, PATTERN, Peptides-struct, and PascalVOC-SP, while ranking second in CLUSTER. Notably, GRASS achieves 21.9% lower MAE in ZINC compared to GRIT, the second-best model on the dataset, which has $O(|V|^{2})$ runtime.
Table 1: Performance on GNN Benchmark Datasets. The performance of GRASS shown here is the mean $\pm$ s.d. of 8 runs on ZINC, and 6 runs on other datasets. The best and second-best results are highlighted. Performance numbers other than that of GRASS are adapted from Ma et al. [47] and Shirzad et al. [61].
Model ZINC MNIST CIFAR10 PATTERN CLUSTER MAE $\downarrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ GCN 0.367 $\pm$ 0.011 90.705 $\pm$ 0.218 55.710 $\pm$ 0.381 71.892 $\pm$ 0.334 68.498 $\pm$ 0.976 GIN 0.526 $\pm$ 0.051 96.485 $\pm$ 0.252 $5$$\pm$ 1.527 85.387 $\pm$ 0.136 64.716 $\pm$ 1.553 GAT 0.384 $\pm$ 0.007 95.535 $\pm$ 0.205 64.223 $\pm$ 0.455 78.271 $\pm$ 0.186 70.587 $\pm$ 0.447 GatedGCN 0.282 $\pm$ 0.015 97.340 $\pm$ 0.143 67.312 $\pm$ 0.311 85.568 $\pm$ 0.088 73.840 $\pm$ 0.326 GatedGCN-LSPE 0.090 $\pm$ 0.001 - - - - PNA 0.188 $\pm$ 0.004 97.94 $\pm$ 0.12 70.35 $\pm$ 0.63 - - DGN 0.168 $\pm$ 0.003 - 72.838 $\pm$ 0.417 86.680 $\pm$ 0.034 - GSN 0.101 $\pm$ 0.010 - - - - CIN 0.079 $\pm$ 0.006 - - - - CRaW1 0.085 $\pm$ 0.004 97.944 $\pm$ 0.050 69.013 $\pm$ 0.259 - - GIN-AK+ 0.080 $\pm$ 0.001 - 72.19 $\pm$ 0.13 86.850 $\pm$ 0.057 - SAN 0.139 $\pm$ 0.006 - - - 76.691 $\pm$ 0.65 Graphormer 0.122 $\pm$ 0.006 - - - - K-Subgraph SAT 0.094 $\pm$ 0.008 - - 86.848 $\pm$ 0.037 77.856 $\pm$ 0.104 EGT 0.108 $\pm$ 0.009 98.173 $\pm$ 0.087 68.702 $\pm$ 0.409 86.821 $\pm$ 0.020 79.232 $\pm$ 0.348 Graphormer-URPE 0.086 $\pm$ 0.007 - - - - Graphormer-GD 0.081 $\pm$ 0.009 - - - - GPS 0.070 $\pm$ 0.004 - 72.298 $\pm$ 0.356 86.685 $\pm$ 0.059 78.016 $\pm$ 0.180 Exphormer - 98.55 $\pm$ 0.039 74.69 $\pm$ 0.125 86.74 $\pm$ 0.015 78.07 $\pm$ 0.037 GRIT 0.059 $\pm$ 0.002 98.108 $\pm$ 0.111 76.468 $\pm$ 0.881 87.196 $\pm$ 0.076 80.026 $\pm$ 0.277 GRASS (ours) 0.046 $\pm$ 0.001 98.905 $\pm$ 0.034 82.471 $\pm$ 0.319 90.997 $\pm$ 0.225 79.347 $\pm$ 0.087
Table 2: Performance on LRGB datasets. The performance of GRASS shown here is the mean $\pm$ s.d. of 8 runs. The best and second-best results are highlighted. Performance numbers other than that of GRASS are adapted from Gutteridge et al. [32], Tönshoff et al. [63], Ma et al. [47], and Shirzad et al. [61]. *Tönshoff et al. [63] has improved the performance of these models via hyperparameter tuning.
Model Peptides-func Peptides-struct PascalVOC-SP AP $\uparrow$ MAE $\downarrow$ Macro F1 $\uparrow$ GCN* 0.6860 $\pm$ 0.0050 0.2460 $\pm$ 0.0007 0.2078 $\pm$ 0.0031 GINE* 0.6621 $\pm$ 0.0067 0.2473 $\pm$ 0.0017 0.2718 $\pm$ 0.0054 GatedGCN* 0.6765 $\pm$ 0.0047 0.2477 $\pm$ 0.0009 0.3880 $\pm$ 0.0040 DIGL+MPNN 0.6469 $\pm$ 0.0019 0.3173 $\pm$ 0.0007 0.2824 $\pm$ 0.0039 DIGL+MPNN+LapPE 0.6830 $\pm$ 0.0026 0.2616 $\pm$ 0.0018 0.2921 $\pm$ 0.0038 MixHop-GCN 0.6592 $\pm$ 0.0036 0.2921 $\pm$ 0.0023 0.2506 $\pm$ 0.0133 MixHop-GCN+LapPE 0.6843 $\pm$ 0.0049 0.2614 $\pm$ 0.0023 0.2218 $\pm$ 0.0174 DRew-GCN 0.6996 $\pm$ 0.0076 0.2781 $\pm$ 0.0028 0.1848 $\pm$ 0.0107 DRew-GCN+LapPE 0.7150 $\pm$ 0.0044 0.2536 $\pm$ 0.0015 0.1851 $\pm$ 0.0092 DRew-GIN 0.6940 $\pm$ 0.0074 0.2799 $\pm$ 0.0016 0.2719 $\pm$ 0.0043 DRew-GIN+LapPE 0.7126 $\pm$ 0.0045 0.2606 $\pm$ 0.0014 0.2692 $\pm$ 0.0059 DRew-GatedGCN 0.6733 $\pm$ 0.0094 0.2699 $\pm$ 0.0018 0.3214 $\pm$ 0.0021 DRew-GatedGCN+LapPE 0.6977 $\pm$ 0.0026 0.2539 $\pm$ 0.0007 0.3314 $\pm$ 0.0024 Transformer+LapPE 0.6326 $\pm$ 0.0126 0.2529 $\pm$ 0.0016 0.2694 $\pm$ 0.0098 SAN+LapPE 0.6384 $\pm$ 0.0121 0.2683 $\pm$ 0.0043 0.3230 $\pm$ 0.0039 GPS+LapPE* 0.6534 $\pm$ 0.0091 0.2509 $\pm$ 0.0014 0.4440 $\pm$ 0.0065 Exphormer 0.6527 $\pm$ 0.0043 0.2481 $\pm$ 0.0007 0.3975 $\pm$ 0.0037 GRIT 0.6988 $\pm$ 0.0082 0.2460 $\pm$ 0.0012 - GRASS (ours) 0.6353 $\pm$ 0.0056 0.2447 $\pm$ 0.0009 0.5633 $\pm$ 0.0092
It has been suggested that oversmoothing can be alleviated by having a shallow GNN, and oversquashing can be avoided by having wide hidden layers [64]. However, narrow-and-deep GNNs can have better expressivity than wide-and-shallow GNNs if the pitfalls of training deep GNNs, such as oversmoothing and oversquashing, can be avoided [43, 44]. This could potentially explain the competitive performance of GRASS which we configured to be narrow and deep in our experiments, as specified in Appendix E. In addition, the datasets we tested are relatively small in sample size (<100k samples), which makes a good inductive bias more important [6, 47], highlighting the ability of GRASS to provide a strong inductive bias on graph-structured data while remaining computationally efficient.
Ablation Study.
GRASS achieves its design goals with multiple components, whose importance is evaluated in the ablation studies shown in Table 3. In particular, the performance of GRASS with different values of $r$, the degree of the superimposed random regular graph, is compared in Table 4.
5 Conclusion
GRASS demonstrates the effectiveness of random rewiring and our graph-tailored attention mechanism by achieving state-of-the-art performance, surpassing more elaborate methods of graph rewiring, as well as more computationally demanding Graph Transformers. To improve the reproducibility of our results, our codebase will be made publicly available on GitHub.
Limitations.
The output of GRASS is inherently random due to random rewiring. The relationship between performance variance and $r$ is demonstrated in Table 4. In scenarios that strictly require deterministic model output, the random number generator used for random rewiring should be seeded with the input graph, making the deployment of GRASS more complicated in these scenarios.
Appendix A Pseudocode of the Permutation Model
1:procedure PermutationModel($r$, $|V|$)
2:     $\sigma\leftarrow\text{2D array of size }(r,|V|)$
3:     for $i\leftarrow 0\text{ to }r-1$ do
4:         $\sigma[i,:]\leftarrow\textsc{RandPerm}(|V|)$triangleright Random permutation of integers between 00 and $|V|-1$
5:     end for
6:     $A\leftarrow\text{Array of size }r\times|V|$triangleright Create an empty adjacency list
7:     for $j\leftarrow 0\text{ to }|V|$ do
8:         for $k\leftarrow 0\text{ to }r$ do
9:              $A[j\times r+k]\leftarrow\{j,\sigma[k,j]\}$triangleright Add an edge to the adjacency list
10:         end for
11:     end for
12:     $A\leftarrow\textsc{RemoveSelfLoop}(A)$triangleright Remove self-loops from the adjacency list
13:     $A\leftarrow\textsc{RemoveMultiEdge}(A)$triangleright Remove multi-edges from the adjacency list
14:     return $A$
15:end procedure
Algorithm 1 The Permutation Model
Appendix B Graphical Descriptions of GRASS
Figure 2: The high-level structure of GRASS.
Figure 3: The structure of an attention layer of GRASS. Node aggregation is attentive, with attention weights derived from edge representations. Edge aggregation is done through an FFN. For simplicity, biases are not shown here.
Appendix C Ablation Study
Table 3: Ablation study results on different components of GRASS. This table shows the performance of each ablated model as the mean $\pm$ s.d. over 8 runs. *The absolute difference between the LapPEs (8 eigenvalues) of neighboring nodes are used as edge encodings, as proposed by Kreuzer et al. [40].
Model ZINC MAE $\downarrow$ GRASS 0.0461 $\pm$ 0.0014       Remove RRWP encoding 0.0797 $\pm$ 0.0040       RRWP encoding $\rightarrow$ relative LapPE* 0.0708 $\pm$ 0.0032       Remove DropKey 0.0549 $\pm$ 0.0125       Remove edge flipping 0.0477 $\pm$ 0.0010       PN-V $\rightarrow$ LN [3] 0.0497 $\pm$ 0.0013       Sum pooling $\rightarrow$ Mean pooling 0.0514 $\pm$ 0.0018
Table 4: Ablation study results on the random regular graph degree $r$. This table shows the performance of GRASS on ZINC with the corresponding values of $r$ as the mean $\pm$ s.d. over 8 runs. The average variance of model performance due to random rewiring, as well as their sample standard deviations, are also shown. They are measured by evaluating the test set 30 times on each trained model. For comparison, the variance of model performance due to randomness in the training process is 2.03e-6 with the hyperparameters shown in Table 7.
Random Regular Graph Degree $r$ 0 3 6 9 12 15 ZINC MAE $\downarrow$ 0.0495 0.0470 0.0461 0.0473 0.0473 0.0478 $\pm$ 0.0016 $\pm$ 0.0023 $\pm$ 0.0014 $\pm$ 0.0015 $\pm$ 0.0024 $\pm$ 0.0011 Variance in ZINC MAE Due to Random Rewiring - 6.31e-8 1.22e-7 1.23e-7 1.79e-7 1.58e-7 - $\pm$ 2.32e-8 $\pm$ 5.90e-8 $\pm$ 9.36e-8 $\pm$ 1.50e-7 $\pm$ 1.04e-7
Appendix D Datasets
Table 5: Statistics of GNN Benchmark Datasets used in section 4, adapted from Rampášek et al. [56].
Dataset # Graphs Avg. # Nodes Avg. # Edges Task Metric ZINC 12000 23.2 24.9 Graph Regression MAE $\downarrow$ MNIST 70000 70.6 564.5 Graph Classif. Accuracy $\uparrow$ CIFAR10 60000 117.6 941.1 Graph Classif. Accuracy $\uparrow$ PATTERN 14000 118.9 3039.3 Node Classif. Accuracy $\uparrow$ CLUSTER 12000 117.2 2150.9 Node Classif. Accuracy $\uparrow$
Table 6: Statistics of LRGB datasets used in section 4, adapted from Dwivedi et al. [24].
Dataset # Graphs Avg. # Nodes Avg. # Edges Avg. Short. Path Avg. Diameter Task Metric Peptides-func 15535 150.94 307.30 20.89 $\pm$ 9.79 56.99 $\pm$ 28.72 Graph Classif. AP $\uparrow$ Peptides-struct 15535 150.94 307.30 20.89 $\pm$ 9.79 56.99 $\pm$ 28.72 Graph Regression MAE $\downarrow$ PascalVOC-SP 11355 479.40 2710.48 10.74 $\pm$ 0.51 27.62 $\pm$ 2.13 Node Classif. Macro F1 $\uparrow$
Appendix E Hyperparameters
Table 7: Model hyperparameters for experiments on GNN Benchmark Datasets. The results are shown in table 1.
Model ZINC MNIST CIFAR10 PATTERN CLUSTER # Parameters 499777 104338 104386 498626 498854 # Attention Layers 49 15 15 53 53 Attention Layer Dim. 32 24 24 32 32 Task Head Hidden Dim. 192 144 144 N/A (Linear) N/A (Linear) # Epochs 2000 200 400 500 100 Warmup Epoch Ratio 0.1 0.05 0.1 0.1 0.1 Batch Size 200 200 200 200 200 Initial Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Peak Learning Rate 5e-4 1e-3 1e-3 1e-3 1e-3 Final Learning Rate 1e-7 1e-7 1e-7 3e-4 1e-7 Betas (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) Weight Decay Factor 0.3 0.3 0.3 3.0 0.3 Attention DropKey Rate 0.1 0.1 0.1 0.3 0.3 Label Smoothing Factor N/A (Regression) 0.1 0.1 0.1 0.1 Random Regular Graph Degree 6 3 3 15 15 RRWP Random Walk Length 32 24 24 32 32
Table 8: Model hyperparameters for experiments on LRGB datasets. The results are shown in table 2.
Model Peptides-func Peptides-struct PascalVOC-SP # Parameters 497866 498155 498133 # Attention Layers 48 48 53 Attention Layer Dim. 32 32 32 Task Head Hidden Dim. 192 192 N/A (Linear) # Epochs 500 500 600 Warmup Epoch Ratio 0.1 0.1 0.1 Batch Size 200 200 200 Initial Learning Rate 1e-7 1e-7 1e-7 Peak Learning Rate 1e-3 1e-3 1e-3 Final Learning Rate 1e-7 1e-7 1e-7 Betas (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) Weight Decay Factor 3.0 3.0 1.0 Attention DropKey Rate 0.1 0.1 0.1 Label Smoothing Factor 0.1 N/A (Regression) 0.1 Random Regular Graph Degree 6 6 3 RRWP Random Walk Length 32 32 16
Appendix F Computational Performance
Table 9: Computational performance of GRASS on GNN Benchmark Datasets. Training time per epoch is the wall-clock time taken to complete a single training epoch, shown as the mean $\pm$ s.d. over all epochs of a single run. Preprocessing time is the wall-clock time taken to load, preprocess, and store the whole dataset prior to training, shown as the mean $\pm$ s.d. over 4 runs. Specifications of the hardware used to run these experiments are also shown.
Dataset ZINC MNIST CIFAR10 PATTERN CLUSTER Training Time per Epoch (s) 2.71 $\pm$ 0.13 11.41 $\pm$ 0.15 13.08 $\pm$ 0.10 150.30 $\pm$ 0.17 122.31 $\pm$ 0.08 Preprocessing Time (s) 28.35 $\pm$ 0.19 74.12 $\pm$ 0.44 139.50 $\pm$ 6.55 33.08 $\pm$ 0.40 27.06 $\pm$ 0.10 Model Compilation Yes No Activation Checkpointing No Yes GPU NVIDIA GeForce RTX 4090 VRAM 24GB CPU AMD Ryzen 9 7950X RAM 128GB
Table 10: Computational performance of GRASS on LRGB datasets. The definitions of these statistics and the procedure of measuring them are identical to that described in Table 9.
Dataset Peptides-func Peptides-struct PascalVOC-SP Training Time per Epoch (s) 10.44 $\pm$ 0.10 10.21 $\pm$ 0.09 21.56 $\pm$ 0.10 Preprocessing Time (s) 69.78 $\pm$ 4.09 69.81 $\pm$ 2.18 181.28 $\pm$ 1.33 Model Compilation Yes Activation Checkpointing No GPU NVIDIA GeForce RTX 4090 NVIDIA RTX A6000 Ada VRAM 24GB 48GB CPU AMD Ryzen 9 7950X RAM 128GB